A simple pendulum of length l and mass m is suspended in a car that is moving with constant velocity v around a circle of radius r. Find its time period and equilibrium position?pls post fast

Grade:12

1 Answers

vikas askiitian expert
13 years ago

car is moving along the circle so centripital force is acting on the car but bob experiences a centrifugal force ....

let the angle made with verticle is @ & T is the tension in the string  then

Tcos@ = mg

Tsin@ = mv2/r

tan@ = v2/rg  

bob is at equilibrium at @ = tan-1 (v2/g)

now

       net force acting on bob = (mv2r) i + (mg) j

                                    a = v2/r i+ g j

               magnitude of a= [ (v2/r)2 + g2 ]

T = 2pi(L/geff)

  =2pi { L/[(v2/r)2 +g2] }1/2